OpenStudy (anonymous):

Let sin p = .71. Which of the following is cos(2p)? A. –.0082 B. .082 C. –.82 D. .0082

jimthompson5910 (jim_thompson5910):

Hint: cos(2p) = 1 - (sin(p))^2

OpenStudy (anonymous):

\[\bf \cos(2x)=1-2\sin^2(x)\]

OpenStudy (anonymous):

@jim_thompson5910 WE THINK ALIKE.

jimthompson5910 (jim_thompson5910):

oh right, there's a 2

jimthompson5910 (jim_thompson5910):

Hint: cos(2p) = 1 - 2*(sin(p))^2 almost alike lol

jimthompson5910 (jim_thompson5910):

this identity (along with others) was found here: http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

@jim_thompson5910 I must congratulate you on your brave efforts.

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

I got 1.50, which clearly isn't an available answer. @jim_thompson5910 @genius12

jimthompson5910 (jim_thompson5910):

you typed it in wrong

jimthompson5910 (jim_thompson5910):

cos(2p) = 1 - 2*(sin(p))^2 cos(2p) = 1 - 2*(0.71)^2 cos(2p) = ???

OpenStudy (anonymous):

A :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct